OpenStudy (anonymous):

Suppose a computer virus begins by infecting 8 computers in the first hour after it is released. Each hour after that, each newly infected computer causes 8 more computers to become infected. The function y = 8x models this situation. Make a table with integer values of x from 1 to 4.

OpenStudy (anonymous):

x=1,2,3,4 y=8,16,24,32

OpenStudy (anonymous):

shouldn't the function be 8^x tho?

OpenStudy (anonymous):

first hour : 8 computers Second hour : 8 * 8 computers = 64 computers, Third : 8 * 64 computers = 512 computers...

OpenStudy (experimentx):

i think it should be rather like this 8, 8 + 64, 72 +(72)*8, 72 +(72)*8 + (72 +(72)*8)*8, ..

OpenStudy (anonymous):

ah, yes, even better. What would be the function of that tho?

OpenStudy (anonymous):

yes your theory is correct to the real world, but the given questions states the relation y=8x

OpenStudy (experimentx):

let's try to find out

OpenStudy (experimentx):

f(x) = 8 * 9^(x-1)

OpenStudy (anonymous):

I only remember the general form of (1+a)^x = y... is it 8(1+8)^(x-1)?

OpenStudy (anonymous):

Excellent! But that doesn't really help the original poster. @Shada, could you confirm that it's indeed 8x or 8^x?

OpenStudy (anonymous):

thanks for trying to help me. i really appreciate it but i dont kno which answer to choose

OpenStudy (anonymous):

The function in your problem, is it : \[8^{x}\] or 8x?

OpenStudy (anonymous):

its 8x

OpenStudy (anonymous):

alright, then Chemengineer was right in the very first answer

OpenStudy (anonymous):

thanks for the help